# Non-Hermitian sensing in the absence of exceptional points

### v1

@@ Abstract
Introduction.—High-precision sensing plays an increasingly important role in modern science, and much effort has been invested in designing novel sensing schemes and next-generation sensors T1986 ; DRC17 . A promising route is to exploit the unique sensitivity of open systems to external perturbations, of which an outstanding example is exceptional-point (EP)-based sensing W14 ; W16 ; LZO+16 ; RHH+17 ; S17 ; L18 ; MGK+18 ; CJL19 ; ZSH+19 ; DPD19 ; MA19 ; W20 ; W20PR ; BB20 ; W20PRA ; WZL+21 ; KS22 ; ASF23 ; WL23 ; Ashida2020 ; Luo2022 ; Bao2021 . While the dynamics of open systems can be described by non-Hermitian effective Hamiltonians, their complex eigenspectra exhibit exotic degeneracies at the branch-point singularities known as the EPs B04 ; H04 ; H12 . A system can exhibit fascinating properties near an EP, where, importantly, the eigenspectra become highly susceptible to weak perturbations, a feature that has been extensively researched for EP-enhanced sensing. Over the past decades, proof-of-principle EP-enhanced sensors have been demonstrated using classical electromagnetic or acoustic waves FSA15 ; HHW+17 ; CKS+17 ; ZPO+18 ; LLS+19 ; XLK+19 ; NAJ+19 ; PNC+20 ; KCE+22 ; JZL+22 ; LLX+23 , and in quantum open systems of trapped ions DSZ+21 , solid spins WLG+19 , or single photons YMT+20 . These prototype sensors and their underlying protocols vitally depend on the presence of EPs, necessitating specific designs of non-Hermitian effective Hamiltonians with fine-tuned parameters.
In this work, we propose and demonstrate non-Hermitian sensing in the absence of EPs. The scheme relies on the non-Hermiticity-enhanced susceptibility of physical observables to weak perturbations, which, distinct from the EP-based sensitivity, is a generic feature of non-Hermitian systems, making the sensing scheme universal CLL+20 . Using a cyclic photonic interferometer, we experimentally simulate the discrete-time dynamics of the probe-field system with tunable non-Hermiticity. We observe enhanced sensitivity under a non-Hermitian setting, thus confirming the feasibility of our scheme. To provide a prototype scenario for application, we then demonstrate the sensing of signals encoded in the setting angle of a wave plate. We also theoretically analyze and experimentally confirm the superiority of the non-Hermitian sensing scheme—it outperforms its Hermitian counterpart in the presence of background noises that do not average out through repetitive measurements.
@@
Theoretical framework.—We consider a generic non-Hermitian qubit sensor perturbed by a weak external field, with the overall Hamiltonian given by $H=\mathcal{H}+\lambda V$, and $\lambda$ denotes the small parameter to be estimated. In a properly chosen orthonormal basis denoted by $\{|0\rangle,|1\rangle\}$, the Hamiltonian of the bare non-Hermitian sensor can be generally expressed as (up to an arbitrary constant)
@@
with $|a|\leq 1$. Its two eigenstates are given by $|\phi_{+}\rangle\sim|0\rangle+\delta|1\rangle$ and $|\phi_{-}\rangle\sim|0\rangle-(\delta/a)|1\rangle$, respectively, with an associated energy splitting $2\mathcal{E}$. Importantly, our sensing scheme assumes the condition $|\delta|\ll 1$, which is always possible by a proper choice of the basis. Since the requirement of $\mathcal{E}\to 0$ is released, the non-Hermitian sensor under study is independent of the properties of any EPs.
@@
$S=\frac{|\mathcal{D}_{t}|^{2}}{|\delta|^{2}+|\mathcal{D}_{t}|^{2}}.$ (3)
Now, consider the time evolution under $H$, where an additional weak field is present. The parameters of the time-evolved state $\{\delta,a,\mathcal{E}\}$ all become $\lambda$-dependent, so is the normalized population in $|0\rangle$, which we denote as $S(\lambda)$. Its sensitivity to $\lambda$ is captured by the susceptibility $\chi_{\lambda}=\partial_{\lambda}S(\lambda)$, which can be decomposed into
@@
The above population-matching condition (namely $|\mathcal{D}_{t}|\sim|\delta|\ll 1$) is achieved by exploiting the generic non-Hermiticity of the sensor, inherently facilitated by the non-orthogonal eigenstates and complex eigenvalues. For example, the non-orthogonality of eigenstates in a pseudo-Hermitian system (where $\mathcal{E}$ is real) allows $a$ to be of the order unity ($a\approx 1$), which straightforwardly yields $\mathcal{D}_{t}\approx i\cot\mathcal{E}t$ CLL+20 . When the evolution time is close to $\mathcal{E}t\simeq\frac{\pi}{2}$, $|\mathcal{D}_{t}|$ could be on the order of $|\delta|$, satisfying the population-matching condition. More general non-Hermitian sensing configurations beyond the pseudo-Hermitian case can also be established based on the population-matching condition supp .
The observations above form the basis of our non-Hermiticity-enhanced sensing scheme, which is fundamentally different from the previously reported EP-enhanced sensing W14 ; W16 . While the latter typically relies on the sensitivity of eigenstates and eigenvalues close to the EP, our scheme exploits the non-Hermitian sensing region illustrated in Fig. 1, which is emergent from the non-Hermitian dynamics starting from a specific initial state in the close proximity to an eigenstate of the unperturbed system.
However, the above discussion does not hold for a Hermitian system, where a unitary time evolution gives $2\mathcal{D}_{t}\approx 1+i\cot\mathcal{E}t$ (where $a=\delta^{2}\ll 1$), and thus $2|\mathcal{D}_{t}|\geq 1$. More specifically, we consider an exemplary Hermitian sensor, where the full Hamiltonian is given by $H_{s}=\lambda\sigma_{x}$. The time-evolved state follows the form of Eq. (2), by taking $\delta=1$ and $\mathcal{D}_{t}=-i\cot\lambda t$, corresponding to the smooth non-sensitive region in Fig. 1. The normalized population in state $|0\rangle$ is
$S(\lambda)=\frac{1}{2}\left(1+\cos 2\lambda t\right).$ (5)
@@
Experiment implementation.—As a paradigmatic example, we experimentally implement the pseudo-Hermitian sensor described above using linear optics, dividing the process into many segments, each of which simulates a fixed time evolution governed by a pseudo-Hermitian Hamiltonian. The process can be extended to a real-time evolution and is regarded as a genuine quantum simulation in linear optics XHS+14 .
In our experiment, we encode the qubit basis in the horizontal and vertical polarization states of photons, i.e., $|0\rangle\equiv|H\rangle$ and $|1\rangle\equiv|V\rangle$. As illustrated in Fig. 2, our experimental setup includes three stages: state initialization, time evolution, and projective measurement. The wave packets of photons are generated by a pulsed laser source with a central wavelength of $808$nm, a pulse width of $88$ps, and a repetition rate of $31.25$kHz. The pulses are attenuated to the single-photon level using neutral density filters. The photons are projected into the initial state |$\ket{H}$ via a polarizing beam splitter (PBS) and then coupled in and out of an interferometric network through a low-reflectivity beam splitter (BS, reflectivity $10\%$), for time evolutions under the non-Hermitian Hamiltonian. Each segment of the time evolution is realized by the interferometric network involving several wave plates and two beam displacers (BDs). Such an interferometric network is capable of realizing an arbitrary $2\times 2$ non-unitary operator TRS18 ; XWZ+19 ; WXB+21 .
@@
$u(\tau)=e^{-i(H-\mathcal{E}\mathbb{I})\tau}=\begin{pmatrix}{\cos\alpha\tau}&{-i\frac{\beta}{\alpha\delta}\sin\alpha\tau}\\ {-i\frac{\beta}{\alpha}\sin\alpha\tau}&{\cos\alpha\tau}\end{pmatrix},$ (6)
@@
Finally, after photons have completed multiple cycles and are coupled out of the network by the BS, they register clicks at an avalanche photo-diode (APD) with a time jitter of $350$ps for detection. The population of photons is obtained through projective measurement, which consists of a PBS and APDs. The counts of the horizontally polarized photons $N_{H}$ and vertically polarized ones $N_{V}$ are registered in the measurement stage and are used to determine susceptibilities in our experiment.
@@
The populations of the state $|H\rangle$ after $5$ segments of the time evolution governed by non-Hermitian and Hermitian Hamiltonians with the initial state $|H\rangle$ are shown in Figs. 3(a) and  3(c), respectively. In the non-Hermitian case, the population changes more sharply compared to its Hermitian counterpart, with respect to small variations in $\lambda$, especially when $\lambda/\varepsilon\in\left[-1.402,1.832\right]$.
@@
Figure 3: Experimental results of the susceptibility of non-Hermitian and Hermitian sensors. The normalized population $S(\lambda)$ of the state $|H\rangle$ and the susceptibility $\chi_{\lambda}$ as a function of $\lambda$ for the non-Hermitian sensor (a)-(b) and its Hermitian counterpart (c)-(d), respectively. The normalized population $S(\theta_{1})$ and the susceptibility $\chi_{\theta_{1}}$ as a function of $\theta_{1}$ for the non-Hermitian sensor (e)-(f) and its Hermitian counterpart (g)-(h), respectively. Symbols represent the experimental data and lines are the corresponding theoretical predictions with the initial state $|H\rangle$ and the evolution time $t=\pi/2\mathcal{E}$. Other parameters are $\delta=0.3015$, $\mathcal{E}=\delta/50$ and $\varepsilon=0.001$. Error bars are due to the statistical uncertainty in photon number counting.
@@
In Figs. 3(b) and  3(d), we show the measured susceptibility $\chi_{\lambda}$ as a function of $\lambda/\varepsilon$ for both the non-Hermitian and Hermitian sensors. Since we take discretized values of $\lambda$ in our experiment, the susceptibility is numerically estimated according to
$\chi_{\lambda_{i}}=\frac{\bar{S}(\lambda_{i+1})-\bar{S}(\lambda_{i})}{\lambda_{i+1}-\lambda_{i}},$ (9)
@@
where the subscript $i$ represents the data index. As illustrated in Fig. 3(b), for the non-Hermitian sensor, we have the measured susceptibility $|\chi_{\lambda}|\leq 1018\pm 97$. Whereas for the Hermitian case in Fig. 3(d), the measured susceptibility $|\chi_{\lambda}|\leq 262\pm 9$. Therefore, the non-Hermitian sensor demonstrates a remarkable advantage for the estimation of $\lambda$.
[…]
@@
For comparison, we use a Hermitian sensor for the same task. We remove the BDs along with $H_{2}$, $H_{H}$, and $H_{V}$, leaving wave plates $Q_{2}(\phi_{2})$-$H_{1}(\theta_{1})$-$Q_{1}(\phi_{1})$. We fix the setting angles of the wave plates to be the same as those of the non-Hermitian sensor, with $\varphi_{1}=0.5\pi$, $\varphi_{2}=-0.5\pi$. Figures 3(g) and  3(h) show the population and the susceptibility of the Hermitian sensor. The measured susceptibility is $|\chi_{\theta_{1}}|\leq 10.8\pm 0.4$. Hence, the non-Hermitian enhancement ratio is $\sim 10.6$.
Non-Hermitian sensing under noise.—Noise is a fundamental challenge for sensors as the sensitivity of a sensor is ultimately limited by noise. Generally, the influence of measurement noises on the sensitivity of the non-Hermitian sensor can be eliminated by repeating the measurement. However, for systematic noises that cannot be averaged out, the non-Hermitian sensors demonstrated here also outperform the Hermitian ones. In the following, we take background noises KG14 ; XZX+21 as an example, and experimentally confirm the non-Hermitian advantage.
We characterize the measurement sensitivity $\Delta S$ using the standard deviation of the parameter to be estimated. The measurement sensitivity indicates the minimum change of the parameter that can be distinguished in the presence of the background noise. This allows us to directly compare the non-Hermitian sensing protocols to more established methods. For simplicity, we focus on the background noise due to the imperfection of detectors. The measured population of the state |$\ket{H}$ in Eq. (8) is then modified as
$\bar{S}^{\prime}=$ $\frac{N_{H}+N^{\prime}_{H}}{N_{H}+N^{\prime}_{H}+N_{V}+N^{\prime}_{V}}$
$\approx$ $\bar{S}+(1-\bar{S})\frac{N_{H}^{\prime}}{N}-\bar{S}\frac{N_{V}^{\prime}}{N}.$ (10)
Here $N^{\prime}_{H,V}$ represents the additional photon numbers in the corresponding polarization states caused by the background noise, and $N=N_{H}+N_{V}$. In general, the background noise is completely random. The corresponding photon numbers are therefore given by $N_{H}^{\prime}\in[0,\eta_{H}N]$ and $N_{V}^{\prime}\in[0,\eta_{V}N]$, respectively. Applying error propagation on Eq. (10), the standard deviation of $\bar{S}^{\prime}$ is
$\Delta\bar{S}^{\prime 2}=$ $(\frac{\partial\bar{S}^{\prime}}{\partial\bar{S}}\Delta\bar{S})^{2}+(\frac{\partial\bar{S}^{\prime}}{\partial N_{H}^{\prime}/N}\Delta\frac{N_{H}^{\prime}}{N})^{2}+(\frac{\partial\bar{S}^{\prime}}{\partial N_{V}^{\prime}/N}\Delta\frac{N_{V}^{\prime}}{N})^{2}$
$=$ $(1-\frac{\eta_{H}+\eta_{V}}{2})^{2}\Delta\bar{S}^{2}+\frac{1}{12}[(1-\bar{S})^{2}\eta_{H}^{2}+\bar{S}^{2}\eta_{V}^{2}].$ (11)
@@
The first term on the right-hand side of Eq. (S6) (in the second line) comes from the fluctuation of photons, which can be averaged out through repetitive measurements. Namely, the term tends to zero when $N\gg 1$. By contrast, the second term comes from the background noise, which is independent of $N$ and cannot be averaged out.
@@
We experimentally demonstrate the performance of the non-Hermitian sensor in the presence of background noise. We use two light-emitting diode (LED) which are placed close to the APDs, to simulate the background noise. We adjust the intensity of the LEDs to ensure that the changes in $N^{\prime}_{H,V}$ are proportional to $N$. For simplicity, we choose $\eta_{V}=1.2\eta_{H}$. We choose the working point of the sensor, i.e., $\lambda$ corresponding to $|\chi_{\lambda}|_{\text{max}}$, and repeat measurements for $50$ times for each of the noise strengths chosen in our experiment. For the measured population of the state |$\ket{H}$, we calculate the standard deviation $\Delta\bar{S}^{\prime}$ based on Eq. (S6) and $\Delta\lambda=\Delta\bar{S}^{\prime}/|\chi_{\lambda}|_{\text{max}}$ through the propagation of error.
In Fig. 4, we show how the measurement uncertainty varies as a function of the noise strength $\eta_{H}$, for both the non-Hermitian and Hermitian sensors. The results show that, with increasing noise strength, the measurement sensitivity $\Delta\lambda$ ($\Delta\theta_{1}$) deteriorates in both cases. However, $\Delta\lambda$ ($\Delta\theta_{1}$) of the non-Hermitian sensor is always smaller compared to the Hermitian one. Thus, the non-Hermitian sensor possesses enhanced sensitivity compared to the Hermitian counterpart in the presence of background noise. As such, non-Hermiticity allows one to enhance the measurement sensitivity, over a conventional Hermitian sensor, both with and without noise.
Conclusion.—We propose and demonstrate a generic non-Hermitian sensing scheme that does not rely on the presence of EPs. Simulating the non-Hermitian dynamics of the sensor-field system, we show that the non-Hermitian sensor is superior to its Hermitian counterpart, both in terms of susceptibility, and in terms of performance against background noise. While the non-Hermitian sensor in our experiment is time-independent, it is possible to achieve time dependence through electro-optical modulators, which should open up fresh opportunities for novel sensor design. Furthermore, while our demonstration is performed on the single-qubit level, it is possible to extend the sensing scheme to multi-qubit cases. Our work thus paves the way toward a new class of non-Hermitian sensors.

### v2

@@ Abstract
Introduction.—High-precision sensing plays an increasingly important role in modern science, and much effort has been invested in designing novel sensing schemes and next-generation sensors [1, 2]. A promising route is to exploit the unique sensitivity of open systems to external perturbations, of which an outstanding example is exceptional-point (EP)-based sensing [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. While the dynamics of open systems can be described by non-Hermitian effective Hamiltonians, their complex eigenspectra exhibit exotic degeneracies at the branch-point singularities known as the EPs [25, 26, 27]. A system can exhibit fascinating properties near an EP, where, importantly, the eigenspectra become highly susceptible to weak perturbations, a feature that has been extensively researched for EP-enhanced sensing. Over the past decades, proof-of-principle EP-enhanced sensors have been demonstrated using classical electromagnetic or acoustic waves [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38], and in quantum open systems of trapped ions [39], solid spins [40], or single photons [41]. These prototype sensors and their underlying protocols vitally depend on the presence of EPs, necessitating specific designs of non-Hermitian effective Hamiltonians with fine-tuned parameters.
In this work, we propose and demonstrate non-Hermitian sensing in the absence of EPs. The scheme relies on the non-Hermiticity-enhanced susceptibility of physical observables to weak perturbations, which, distinct from the EP-based sensitivity, is a generic feature of non-Hermitian systems, making the sensing scheme universal [42, 43, 44]. Using a cyclic photonic interferometer, we experimentally simulate the discrete-time dynamics of the probe-field system with tunable non-Hermiticity. We observe enhanced sensitivity under a non-Hermitian setting, thus confirming the feasibility of our scheme. To provide a prototype scenario for application, we then demonstrate the sensing of signals encoded in the setting angle of a wave plate. We also theoretically analyze and experimentally confirm the superiority of the non-Hermitian sensing scheme—it outperforms its Hermitian counterpart in the presence of background noises that do not average out through repetitive measurements.
@@
Theoretical framework.—We consider a generic non-Hermitian qubit sensor perturbed by a weak external field, with the overall Hamiltonian given by $H=\mathcal{H}+\lambda V$, and $\lambda$ denotes the small parameter to be estimated and $V=\sigma_{x}$. In a properly chosen orthonormal basis denoted by $\{|0\rangle,|1\rangle\}$, the Hamiltonian of the bare non-Hermitian sensor can be generally expressed as (up to an arbitrary constant)
@@
with $|a|\leq 1$. Its two eigenstates are given by $|\phi_{+}\rangle\sim|0\rangle+\delta|1\rangle$ and $|\phi_{-}\rangle\sim|0\rangle-(\delta/a)|1\rangle$, respectively, with an associated energy splitting $2\mathcal{E}$. Importantly, our sensing scheme assumes the condition $|\delta|\ll 1$, which is always possible by a proper choice of the basis. Since the requirement of $|\mathcal{E}|\to 0$ is relaxed, the non-Hermitian sensor under study is independent of the properties of any EPs. The working parameters selected for our experiments below are all far away from any EP, as characterized by the Frobenius norm distance [45, 46].
@@
$S(t)=\frac{|\mathcal{D}_{t}|^{2}}{|\delta|^{2}+|\mathcal{D}_{t}|^{2}}.$ (3)
Now, consider the time evolution under $H$, where an additional weak field is present. The parameters of the time-evolved state $\{\delta,a,\mathcal{E}\}$ all become $\lambda$-dependent, so is the normalized population in $|0\rangle$, which we denote as $S(\lambda,t)$. Its sensitivity to $\lambda$ is captured by the susceptibility $\chi_{\lambda}=\partial_{\lambda}S(\lambda,t)$, which can be decomposed into
@@
The above population-matching condition (namely $|\mathcal{D}_{t}|\sim|\delta|\ll 1$) is achieved by exploiting the generic non-Hermiticity of the sensor, inherently facilitated by the non-orthogonal eigenstates and complex eigenvalues. For example, the non-orthogonality of eigenstates in a pseudo-Hermitian system (where $\mathcal{E}$ is real) allows $a$ to be of the order unity ($a\approx 1$), which straightforwardly yields $\mathcal{D}_{t}\approx i\cot\mathcal{E}t$ [42]. When the evolution time is close to $\mathcal{E}t\simeq\frac{\pi}{2}$, $|\mathcal{D}_{t}|$ could be on the order of $|\delta|$, satisfying the population-matching condition. More general non-Hermitian sensing configurations with gain and loss beyond the pseudo-Hermitian case can also be established based on the population-matching condition [46].
The observations above form the basis of our non-Hermiticity-enhanced sensing scheme, which is fundamentally different from the previously reported EP-enhanced sensing [3, 4]. While the latter typically relies on the sensitivity of eigenstates and eigenvalues close to the EP, our scheme exploits the non-Hermitian sensing region illustrated in Fig. 1, which is emergent from the non-Hermitian dynamics starting from a specific initial state in the close proximity to an eigenstate of the unperturbed system.
However, the above discussion does not hold for a Hermitian system, where a unitary time evolution gives $2\mathcal{D}_{t}\approx 1+i\cot\mathcal{E}t$ (where $a=\delta^{2}\ll 1$), and thus $2|\mathcal{D}_{t}|\geq 1$. More specifically, we consider an exemplary Hermitian sensor, where the full Hamiltonian is given by $H_{s}=\lambda\sigma_{x}$. The time-evolved state follows the form of $|\psi(t)\rangle=e^{-i\lambda\sigma_{x}t}|\psi(0)\rangle$. Hence, the normalized population in state $|0\rangle$ is
$S(\lambda,t)=\frac{1}{2}\left(1+\cos 2\lambda t\right).$ (5)
@@
Experiment implementation.—As a paradigmatic example, we experimentally implement the pseudo-Hermitian sensor described above using linear optics, dividing the process into many segments, each of which simulates a fixed time evolution governed by a pseudo-Hermitian Hamiltonian. The process can be extended to a real-time evolution and is regarded as a genuine quantum simulation in linear optics [47].
In our experiment, we encode the qubit basis in the horizontal and vertical polarization states of photons, i.e., $|0\rangle\equiv|H\rangle$ and $|1\rangle\equiv|V\rangle$. As illustrated in Fig. 2, our experimental setup includes three stages: state initialization, time evolution, and projective measurement. The wave packets of photons are generated by a pulsed laser source with a central wavelength of $808$nm, a pulse width of $88$ps, and a repetition rate of $31.25$kHz. The pulses are attenuated to the single-photon level using neutral density filters. The photons are projected into the initial state |$\ket{H}$ via a polarizing beam splitter (PBS) and then coupled in and out of an interferometric network through a low-reflectivity beam splitter (BS, reflectivity $10\%$), for time evolutions under the non-Hermitian Hamiltonian. Each segment of the time evolution is realized by the interferometric network involving several wave plates and two beam displacers (BDs). Such an interferometric network is capable of realizing an arbitrary $2\times 2$ non-unitary operator [48, 49, 50].
@@
$u(\tau)=e^{-i(H-\mathcal{E}\mathbb{I})\tau}=\begin{pmatrix}{\cos\alpha\tau}&{-i\frac{\alpha}{\beta}\sin\alpha\tau}\\ {-i\frac{\beta}{\alpha}\sin\alpha\tau}&{\cos\alpha\tau}\end{pmatrix},$ (6)
@@
Finally, after photons have completed multiple cycles and are coupled out of the network by the BS, they register clicks at an avalanche photo-diode (APD) with a time jitter of $350$ps for detection. The population of photons is obtained through projective measurement, which consists of a PBS and APDs. The counts of the horizontally polarized photons $N_{H}$ and vertically polarized ones $N_{V}$ are registered in the measurement stage and are used to determine susceptibilities in our experiment [46].
@@
The populations of the state $|H\rangle$ after $5$ segments of the time evolution governed by non-Hermitian and Hermitian Hamiltonians with the initial state $|H\rangle$ are shown in Figs. 3(a) and  3(c), respectively. In the non-Hermitian case, the population changes more sharply compared to its Hermitian counterpart, with respect to small variations in $\lambda$, especially when $\lambda\in\left[-0.105,0.138\right]$.
@@
Figure 3: Experimental results of the susceptibility of non-Hermitian and Hermitian sensors. The normalized population $S(\lambda)$ of the state $|H\rangle$ and the susceptibility $\chi_{\lambda}$ as a function of $\lambda$ for the non-Hermitian sensor (a)-(b) and its Hermitian counterpart (c)-(d), respectively. The normalized population $S(\theta_{1})$ and the susceptibility $\chi_{\theta_{1}}$ as a function of $\theta_{1}$ for the non-Hermitian sensor (e)-(f) and its Hermitian counterpart (g)-(h), respectively. Symbols represent the experimental data and lines are the corresponding theoretical predictions with the initial state $|H\rangle$ and the evolution time $t=\pi/2\mathcal{E}$. Other parameters are $\delta=0.3$ and $\mathcal{E}=0.5$. Error bars are due to the statistical uncertainty in photon number counting.
@@
In Figs. 3(b) and  3(d), we show the measured susceptibility $\chi_{\lambda}$ as a function of $\lambda$ for both the non-Hermitian and Hermitian sensors. Since we take discretized values of $\lambda$ in our experiment, the susceptibility is numerically estimated according to
$\chi_{\lambda_{i}}=\frac{S_{i+1}-S_{i}}{\lambda_{i+1}-\lambda_{i}},$ (9)
@@
where the subscript $i$ represents the data index. As illustrated in Fig. 3(b), for the non-Hermitian sensor, we have the measured susceptibility $|\chi_{\lambda}|\leq 19\pm 1$. Whereas for the Hermitian case in Fig. 3(d), the measured susceptibility $|\chi_{\lambda}|\leq 3.8\pm 0.6$. Therefore, the non-Hermitian sensor demonstrates a remarkable advantage for the estimation of $\lambda$.
An application.—In the experiment above, the external field $V$ is simulated by a set of linear optical elements, and the adjustment of the parameter $\lambda$ is realized by tuning the setting angles of the combination of wave plates (H1, H2 and HV). As a primordial demonstration of non-Hermitian sensing, we now encode the signal $\lambda$ in the setting angle of a single wave plate. Our goal is to estimate the parameter $\theta_{1}$ of the wave plate $H_{1}$ with the non-Hermitian sensor. For this purpose, photons are initialized in the state $|\psi(0)\rangle=|H\rangle$. We fix the setting angles of the other wave plates as $\varphi_{1}=0.5\pi$, $\varphi_{2}=-0.5\pi$, $\theta_{2}=0.03\pi$, $\theta_{V}=0.43\pi$, which are determined by numerical simulations and optimization analysis. The populations of the state $|H\rangle$ after $5$ segments of the time evolution governed by the non-Hermitian Hamiltonian with the initial state $|H\rangle$ are shown in Fig. 3(e). Apparently, the measured population becomes more sensitive to $\theta_{1}$ round $\theta_{1}=-0.06\pi$. The susceptibility $\chi_{\theta_{1}}$ is calculated from the measured population and shown in Fig. 3(f). The maximum value of the susceptibility is $115\pm 18$.
@@
For comparison, we use a Hermitian sensor for the same task. We remove the BDs along with $H_{2}$, $H_{H}$, and $H_{V}$, leaving wave plates $Q_{2}(\phi_{2})$-$H_{1}(\theta_{1})$-$Q_{1}(\phi_{1})$. We fix the setting angles of the wave plates to be the same as those of the non-Hermitian sensor, with $\varphi_{1}=0.5\pi$, $\varphi_{2}=-0.5\pi$. Figures 3(g) and  3(h) show the population and the susceptibility of the Hermitian sensor. The measured susceptibility is $|\chi_{\theta_{1}}|\leq 11\pm 2$. Hence, the non-Hermitian enhancement ratio is $\sim 10.5$.
Non-Hermitian sensing under noise.—Noise is a fundamental challenge for sensors as the performance of a sensor is ultimately limited by noise. The non-Hermitian sensors demonstrated here outperform the Hermitian ones, particularly when the technical noises in experiments cannot be averaged out through repetitive measurements. Paradigmatic scenarios are sensing processes with strong background detection noises, or systematic errors; the latter usually results from certain unknown consistent deviations of the real experimental setting from the a priori scenario expected by the measurement agent [46]. In the following, we take background noises [51, 52] as an example, and experimentally confirm the non-Hermitian advantage. For simplicity, we focus on the background noise due to the imperfection of detectors. The measured population of the state |$\ket{H}$ in Eq. (8) is then modified as
$S^{\prime}=$ $\frac{N_{H}+N^{\prime}_{H}}{N_{H}+N^{\prime}_{H}+N_{V}+N^{\prime}_{V}}$
$\approx$ $S+(1-S)\frac{N_{H}^{\prime}}{N}-S\frac{N_{V}^{\prime}}{N}.$ (10)
Here $N^{\prime}_{H,V}$ quantify the additional noisy photons in the corresponding polarization states caused by the background noise, and $N=N_{H}+N_{V}$. In general, the background noise is completely random. The corresponding photon numbers are assumed to be uniformly distributed, with $N_{H}^{\prime}\in[0,\eta_{H}N]$ and $N_{V}^{\prime}\in[0,\eta_{V}N]$, respectively. With saturated photon detectors [53, 54], one can increase $N$ for better measurement precision, by considering a larger interrogation time, which typically involves more background photons as well. Hence, we assume fixed values of $\eta_{H,V}$ in the following. Applying the error propagation procedure on Eq. (10), the measurement noise of $S^{\prime}$ is
$\Delta S^{\prime 2}=$ $(\frac{\partial S^{\prime}}{\partial S}\Delta S)^{2}+(\frac{\partial S^{\prime}}{\partial N_{H}^{\prime}/N}\Delta\frac{N_{H}^{\prime}}{N})^{2}+(\frac{\partial S^{\prime}}{\partial N_{V}^{\prime}/N}\Delta\frac{N_{V}^{\prime}}{N})^{2}$
$=$ $(1-\frac{\eta_{H}+\eta_{V}}{2})^{2}\Delta S^{2}+\frac{1}{12}[(1-S)^{2}\eta_{H}^{2}+S^{2}\eta_{V}^{2}].$ (11)
@@
The first term on the right-hand side of Eq. (11) (in the second line) comes from the fluctuation of photons, which can be averaged out through repetitive measurements. Namely, the term tends to zero when $N\gg 1$. By contrast, the second term comes from the background noise, which is independent of $N$ and cannot be averaged out.
@@
We experimentally demonstrate the performance of the non-Hermitian sensor in the presence of background noise. We use two light-emitting diode (LED) which are placed close to the APDs, to simulate the background noise. We adjust the intensity of the LEDs to ensure that the changes in $N^{\prime}_{H,V}$ are proportional to $N$. For simplicity, we choose $\eta_{V}=1.2\eta_{H}$. We choose the working point of the sensor, i.e., $\lambda$ corresponding to $|\chi_{\lambda}|_{\text{max}}$, and repeat measurements for $50$ times for each of the noise strengths chosen in our experiment. For better contrast, we adopt the same integration time and photon counts for the non-Hermitian and Hermitian sensors, to simulate the saturation for detectors. Although the input photon intensity in the non-Hermitian case is higher than that in the Hermitian case, it should not matter for saturated detectors, as the measurement results would be the same when further increasing the input photon intensity of the Hermitian detector. For the measured population of the state |$\ket{H}$ with the measurement noise $\Delta S^{\prime}$ in Eq. (11), the measurement uncertainty of the parameter $\lambda$, namely the minimum error between the estimated and true values of $\lambda$, is quantified by $\Delta\lambda\approx\Delta S^{\prime}/|\chi_{\lambda}|_{\text{max}}$ through the propagation of error  [46].
In Fig. 4, we show how the measurement uncertainty varies as a function of the noise strength $\eta_{H}$, for both the non-Hermitian and Hermitian sensors. The results show that, with increasing noise strength, the measurement uncertainty $\Delta\lambda$ ($\Delta\theta_{1}$) deteriorates in both cases. However, $\Delta\lambda$ ($\Delta\theta_{1}$) of the non-Hermitian sensor is always smaller compared to the Hermitian one. Thus, the non-Hermitian sensor possesses improved performance compared to the Hermitian counterpart in the presence of background noise. As such, non-Hermiticity allows one to reduce the measurement uncertainty, over a conventional Hermitian sensor, both with and without noise.
Discussions.— We propose and demonstrate a generic non-Hermitian sensing scheme that does not rely on the presence of EPs. Simulating the non-Hermitian dynamics of the sensor-field system, we show that the non-Hermitian sensor is superior to its Hermitian counterpart, both in terms of susceptibility, and in terms of performance against technical noises. While the non-Hermitian sensor in our single-photon experiment is time-independent, it is possible to achieve time dependence through electro-optical modulators, which should open up fresh opportunities for novel sensor design. Our scheme exploiting a general non-Hermitian Hamiltonian is also feasible in other physical systems, such as the solid-state spins [40], superconducting transmon circuits, and single trapped ions [39, 34], where non-Hermitian physics are routinely simulated on the single-qubit basis and beyond. Furthermore, while our demonstration is performed on the single-qubit level, it is possible to extend the sensing scheme to multi-qubit cases. Our work thus paves the way toward a new class of non-Hermitian sensors.